Hello, I was given this question and I have no idea what it means.
Find all real solutions (m)^-4/3 = 16 ??????? Could you please help me with this one. Thank you, Paul

Answer by edjones(8007)

1/m^4/3=16 negative sign in front of exponent means the reciprocal.
We must reverse the process to see what we started with.
1/x =16 substitute x for m^4/3
x=1/16 multiply each side by x/16
1/16^3=1/4096
1/4096^(1/4)=1/8
m=1/8
Check:
(1/8)^-4/3=8^4/3
8^(1/3)=2
2^4=16
Ed